ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 100, 1997

Alon, N., J-H. Kim and J. Spencer Nearly perfect matchings in regular simple hypergraphs

171

Bates, S. M. On smooth, nonlinear surjections of Banach spaces  209

Benjamini, I. On the support of harmonic measure for the random walk

1

Buzzi, J. Intrinsic ergodicity of smooth interval maps  125

Caenepeel, S., G. Militaru and S. Zhu Crossed modules and Doi–Hopf modules

221

Dillen, F., M. Petrovic and L. Verstraelen Einstein, conformally flat and semi-symmetric submanifolds satisfying Chen’s equality

163

Dong, X.H. On a theorem of Bazilevic for areally mean p-valent functions attaining maximal growth on k rays

327

Downey, R. and M. Stob Minimal pairs in initial segments of the recursively enumerable degrees

7

Ferenczi, S. Measure-theoretic complexity of ergodic systems  189

Ginzburg, D., S. Rallis and D. Soudry On the automorphic theta representation for simply laced groups

61

Hamachi, T. Suborbits and group extensions of flows  249

Kim, J-H. See N. Alon

Kinnunen, J. The Hardy–Littlewood maximal function of a Sobolev function

117

Lee, S. T. and C.-B. Zhu Degenerate principal series and local theta correspondence II

29

Militaru, G. See S. Caenepeel

Moran, W. and A. D. Pollington The discrimination theorem for normality to non-integer bases

339

Musson, I. M. Links between cofinite prime ideals in quantum function algebras

285

Nerurkar, M. G. and H. J. Sussmann Construction of minimal cocycles arising from specific differential equations

309

Petrovic, M. See F. Dillen

Pollington, A. D. See W. Moran Rallis, S. See D. Ginzburg Soudry, D. See D. Ginzburg Spencer, J. See N. Alon Stob, M. See R. Downey Sussmann, H. J. See M. G. Nerurkar Verstraelen, L. See F. Dillen Zhu, C.-B. See S. T. Lee Zhu, S. See S. Caenepeel